OpenStudy (anonymous):

Determine the slope of the following linear equation. 7y-2x=0

OpenStudy (jonnyvonny):

The slope of any linear function is the value associated with the x-value. However, we must get y alone so we may see what the function would be. \[7y-2x=0\rightarrow7y=2x rightarrowy=(2/7)x\]

OpenStudy (jonnyvonny):

y=(2/7)x**